OpenStudy (anonymous):

A nurse holding a newborn 3.0-kg baby at a height of 1.2 m off the floor carries the baby 15 m at constant velocity along a hospital corridor. How much work has the force of gravity done on the baby?

OpenStudy (anonymous):

m = 3 kg h = 1.2 m delta d = 15 m g = 9.8 N/kg Fg = ? F = Fg = mg = (3 kg)(9.8 N/kg) = 29.4 N

OpenStudy (amistre64):

how far has the baby moved in the direction of gravity?

OpenStudy (anonymous):

zero.

OpenStudy (amistre64):

then by all accounts; gravity hasnt done any work

OpenStudy (anonymous):

ohh..

OpenStudy (anonymous):

how do I know whether work is zero, and if gravity hasn't done any work?

OpenStudy (amistre64):

gravity moves things in one direction; down. if there is no distance moved doward, no work is done in that direction

OpenStudy (anonymous):

ohhh.. thank you!

OpenStudy (amistre64):

if the nurse lifts the baby; then work is done against gravity

OpenStudy (anonymous):

oh but in this case it was asking how much work has the force of gravity done on the baby..

OpenStudy (anonymous):

does lifting the baby a part of the question?

OpenStudy (amistre64):

lifting is not mention in the question; but say the question says: how much work is done to life the baby up to 1.5 meters then we do the W = F*d stuff

OpenStudy (amistre64):

3 kg * 1.5 m = 4.5 kg/m of work

OpenStudy (anonymous):

ohhh.. okay, thank you! :)